OpenStudy (anonymous):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

OpenStudy (anonymous):

m=y2-y1/x2-x1 m=-3-9/2-2 m=-12/0

OpenStudy (anonymous):

The slope will be 0 :/

OpenStudy (anonymous):

I think it is called "undefined" but i am not sure..

OpenStudy (anonymous):

Thanks @markz123 & @xoya

OpenStudy (anonymous):

Your always welcome :)

OpenStudy (anonymous):

no problem:]